# Erasing Without Remembering: Implicit Knowledge Forgetting in Large Language Models

### v2

@@ Abstract
In this paper, we investigate knowledge forgetting in large language models with a focus on its generalisation—ensuring that models forget not only specific training samples but also related implicit knowledge. To this end, we begin by identifying a broader unlearning scope that includes both target data and logically associated samples, including rephrased, subject-replaced, one-hop reasoned, and relation-reversed data. To rigorously evaluate generalisation, we introduce UGBench—the first comprehensive benchmark specifically designed to assess the unlearning of in-scope implicit knowledge covering 13 state-of-the-art methods across three datasets. UGBench reveals that unlearned models can still recall paraphrased answers and retain target facts in intermediate layers. This motivates us to take a preliminary step toward more generalised implicit knowledge forgetting by proposing PerMU—a novel probability perturbation-based unlearning paradigm. PerMU simulates adversarial unlearning samples to eliminate fact-related tokens from the logit distribution, collectively reducing the probabilities of all answer-associated tokens. Experiments are conducted on a diverse range of datasets, including TOFU, Harry Potter, ZsRE, WMDP, and MUSE, using models ranging from 1.3B to 13B in scale. The results demonstrate that PerMU delivers up to a 50.40% improvement in unlearning vanilla target data while maintaining a 40.73% boost in forgetting implicit knowledge. Our code can be found in https://github.com/MaybeLizzy/UGBench.
@@ 1 Introduction
To further substantiate this observation, our first contribution in this paper is to introduce UGBench, the first comprehensive benchmark specifically designed to evaluate the unlearning ability of existing methods in forgetting neighbouring implicit knowledge. The benchmark spans three data domains: two widely-used machine unlearning datasets, TOFU [18] and the Harry Potter books [19, 20], as well as a popular model editing dataset, ZsRE [21]. Evaluations are performed on 13 existing methods using two language models of different scales, Phi-1.3B and LLaMA2-7B. Our empirical analyses on UGBench uncover the following unique findings that have been overlooked by existing research:
@@
[…]
@@
Motivated by the findings from UGBench, we highlight a critical challenge in LLM unlearning: enhancing generalisation in forgetting implicit knowledge. However, achieving this goal is not without challenges. One vanilla approach is to identify and label all relevant latent knowledge (e.g., rephrased versions and paraphrased answers) for unlearning. Yet, this process is prohibitively labour-intensive and impractical, motivating the development of a novel solution to enhance generalisation.
@@
Building on the analysis using MSM, PerMU injects random noise into the embeddings of the most sensitive tokens, disrupting the model’s ability to recall factual information. As a result, the top-ranked tokens in the next-token probability distribution shift away from fact-related terms and toward tokens driven primarily by grammatical structure or contextual cues. Since the probability distribution reflects a language model’s internal knowledge [22], directly adjusting the logit distribution to suppress fact-related information offers an intuitive and effective unlearning strategy. PerMU implements this by subtracting the original distribution from the perturbed one, then fine-tuning the model to minimize the distance to this residual distribution—ensuring that correct fact-related tokens are assigned lower probabilities, thereby achieving unlearning. As such, PerMU simultaneously addresses both causes of the generalisation challenge identified in UGBench as follows:
[…]
In sum, our contributions are twofold: (1) a novel evaluation benchmark, UGBench, designed to assess the generalisation capability of existing methods in forgetting implicit knowledge; and (2) an unlearning method based on probability perturbation, PerMU, that effectively prevents models from recalling associated facts. We evaluate PerMU across five data domains, including the WMDP [23] and MUSE [24] datasets, using models of varying scales (1.3Bsimilar-$\sim$13B). Comparative experiments show that PerMU achieves up to a 50.40% improvement in unlearning and a 40.73% enhancement in generalisation, all while maintaining high model utility and superior generation quality.
@@ 2 Related Work
We provide a brief overview of existing machine unlearning methods for LLMs, categorised into training-free and training-based approaches [6], along with the evaluation methods relevant to our proposed UGBench benchmark.
@@
Unlearning Evaluation. Several studies have examined LLM unlearning from different perspectives [29, 30, 31, 32, 33]. Specifically, Patil et al. [25] investigate the effectiveness of typical model editing methods in removing information from model weights. Hong et al. [34] use vocabulary projections to analyze concept vectors through parametric knowledge traces. Yao et al. [35] evaluate seven different unlearning methods on longer-context tasks across three source domains. Meanwhile, Shi et al. [24] assess six desirable properties of unlearned models across eight unlearning methods. Jia et al. [20] focus on the influence of second-order optimization on unlearning. Li et al. [23] examine malicious use scenarios in biosecurity, cybersecurity, and chemical security. Qiu et al. [36] evaluate four distinct unlearning methods for removing highly interconnected data. Du et al. [37] investigate the risk of knowledge leakage after unlearning. The survey by Liu et al. [17] highlights often-overlooked aspects of existing LLM unlearning research and introduces the concept of unlearning scope. In contrast, UGBench offers a more precise and formalized definition of unlearning scope, marking a fundamental distinction from prior work. While Patil et al. [31] explore extraction attacks to recover “deleted” information from intermediate hidden states, their focus lies primarily on model editing techniques such as ROME [38]. In comparison, UGBench provides a comprehensive evaluation of the generalisation ability of existing machine unlearning methods across multiple dimensions. Regarding the generalisation evaluation, Dang et al. [32] focuses on improving generalised model utility, by introducing small Gaussian noise to retain data. Unlike prior works, we aim to enhance the generalisation in unlearning implicit knowledge associated with the target fact.
@@ 2.1 Problem Definition
The objective of machine unlearning is to enable an initial target model to forget specific unlearning samples as if it were never trained on them, while preserving the model’s performance on unrelated knowledge. More specifically, the target model $f_{\theta_{tr}}$ is represented by a function $f:\mathbb{X}\mapsto\mathbb{Y}$, where $\theta_{tr}$ denotes the parameters of the target model. Let the pre-training dataset be $D_{tr}$, and the dataset to be forgotten be $D_{f}$. The retained dataset is then defined as $D_{r}=D_{tr}\backslash D_{f}$. The ideal retained model, $f_{\theta_{r}}$, is one that has never been trained on $D_{f}$. Since $\theta_{tr}$ is not directly accessible, we define an unlearning procedure $\mathbb{U}$, which takes $f_{\theta_{tr}}$ and $D_{f}$ as inputs, producing an unlearned model $f_{\theta_{u}}\sim\mathbb{U}(f_{\theta_{tr}},D_{f})$. The unlearned model’s predictions should also change for the paraphrased forget dataset $D_{p}$. Therefore, given a distance metric $m(\cdot)$, the objective of the unlearning algorithm is to minimize the distance between $f_{\theta_{u}}$ and $f_{\theta_{r}}$ for each sample $x\in D_{f}\cup D_{p}$ : $\frac{\mathbb{E}[m(f_{\theta_{u}}(x))]}{\mathbb{E}[m(f_{\theta_{r}}(x))]}\approx 1.$
3 Benchmark Setup
In this section, we provide a detailed overview of the experimental setups for the proposed unlearning generalization benchmark, UGBench, covering the datasets, evaluation metrics, baselines, and implementation details.
@@ 3.1 Dataset
UGBench is evaluated across three distinct data domains, including two widely-used machine unlearning datasets, TOFU [18] and Harry Potter (HP)  [19], as well as a popular model editing dataset, ZsRE [21].
@@ 3.2 Evaluation Metrics
  • Truth Ratio (TR): For a given question, we compute a ratio that approximately compares how likely its correct answer is to an incorrect answer. Let $\hat{a}$ denote a paraphrased version of the correct answer, $\mathcal{A}_{\text{pert }}$ is the set of paraphrased incorrect answer. The truth ratio can be written as:
@@
    We report $\mathrm{TR}=R_{\text{truth }}$ on forget set, and $\mathrm{TR}=max(0,1-R_{\text{truth }})$ on retain set. Therefore, the Truth Ratio score is expected to be higher on both the retain set and the forget set.
@@
3.3 Thirteen Benchmarked Algorithms
We evaluate 13 efficient unlearning methods, including Gradient Ascent (GA[10], Direct Preference Optimization (DPO[41], Negative Preference Optimization (NPO[11], Task Vectors (TV[13], Who’s Harry Potter (WHP[19], ULD [14] and ICL [26, 8].
@@ 3.3 Thirteen Benchmarked Algorithms
We combine GA, DPO, and NPO with these two regularizations using a retain weight $RW$, denoted as “+GDR” or “+KLR”. The total unlearning loss is given by $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{unlearn}}(\theta)+RW\cdot\mathcal{L}_{\mathrm{retain}}(\theta)$. Following previous work [24, 18], we set $RW=1$. The combination of GDR and KLR results in a total of thirteen unlearning methods evaluated in UGBench. Retain refers to the retain model, fine-tuned exclusively on the retain set without exposure to any forget data, and is considered an upper bound.
@@ 3.4 Implementation Details
UGBench is tested across two models, Phi-1.3B and LLaMA2-7B. When tested on TOFU, we use the checkpoints of the pre-trained target model from the TOFU Leaderboard [^1] https://huggingface.co/spaces/locuslab/tofu_leaderboard. For the Harry Potter and ZsRE datasets, we first fine-tune the model on the respective dataset before applying unlearning. The fine-tuning settings are as follows: learning rate of 3e-5, 10 epochs, batch size of 8, with a gradient accumulation step of 4. For Task Vector and WHP, to obtain the reinforced model for unlearning, we fine-tune the target model for 10 epochs using the same learning rate and batch size. For ULD, we obtain the assistant model by fine-tuning the target model using the default settings provided by [14]. For the unlearning process, the unlearning batch size is set to 8, with a gradient accumulation step of 4. The process is conducted over 5 epochs, using a default learning rate of 2e-5. Since different learning rates can result in varying trade-offs between forgetting and retention, we slightly adjust the learning rate for each method to ensure comparable levels of model utility. To ensure fairness, all other unlearning hyper-parameters follow the default settings for each respective unlearning algorithm. All results are averaged over three runs. We use one A100 GPU with 80 GB of RAM. Note that during fine-tuning and unlearning on LLaMA2-7B, we update all 7B model parameters.
4 Benchmark Analysis
In this section, we present a comprehensive benchmark analysis. Through extensive experiments, we observe that existing unlearning methods exhibit limited generalization ability. A detailed explanation is provided below.
@@ 5 Proposed Approach: PerMU
Driven by the insights from UGBench, to achieve our goal of more generalised unlearning, simply reversing the gradient descent loss on forget samples—a common practice—is insufficient to fully erase the memory of implicit knowledge. Recent studies have shown that LM’s internal knowledge is largely captured in its probability distribution [22], which has motivated state-of-the-art unlearning approaches like ULD [14]. These methods suppress fact-related information by modifying the logit distribution: an assistant model is fine-tuned to overfit the ground truth, producing sharp logits that are then subtracted from the original logits to reduce the probability of correct responses.
However, this fine-tuning process narrowly targets ground truth answers while neglecting paraphrased or semantically related alternatives. Consequently, these related tokens may still remain in the subtracted logits, posing inference-time risks of generating semantically equivalent answers. Moreover, these methods are computationally expensive due to additional training and rely heavily on the quality of the assistant model, limiting their scalability and practicality in real-world applications.
Motivated by these limitations, we propose PerMU, a more robust and efficient logit-adjustment method for achieving generalised unlearning. Unlike prior approaches that rely on assistant models, PerMU generates adversarial unlearning samples to jointly suppress all answer-related tokens in the logit space. This is achieved through computationally efficient perturbations that produce a modified distribution, which is then used for logit subtraction and matching. To effectively identify which tokens to perturb, we introduce a theoretically grounded model sensitivity metric, MSM, and present the complete unlearning pipeline in the following subsection.
@@ 5.1 Model Sensitivity Metric
Let the parameters of the target model $\mathrm{LM}$ be $\mathbf{W}\in\mathbb{R}^{n\times n}$, where $n$ represents the hidden dimension of $\mathrm{LM}$. Given an unlearning sample $x$ with a sequence length of $m$, the $i$-th token is represented as $x_{i},i\in\{1,2,\dots,m\}$, and the perturbation applied to this token is denoted by $\Delta_{i}$. Then the perturbed token can be represented as $\hat{x_{i}}=x_{i}+\Delta_{i}$. The change in the model’s output due to $\Delta_{i}$ is measured in terms of the loss $\mathcal{J}(x_{i},\hat{x_{i}})$. The relationship between $\Delta_{i}$ and $\mathcal{J}(x_{i},\hat{x_{i}})$ is not strictly linear. If the model is resilient to certain tokens, even a larger $\Delta_{i}$ results in a relatively small change in $\mathcal{J}(x_{i},\hat{x_{i}})$. Conversely, if the model is sensitive to specific tokens, a small $\Delta_{i}$ can cause a significant difference in $\mathcal{J}(x_{i},\hat{x_{i}})$.
To quantify the extent of perturbation a model can tolerate, we draw inspiration from Zhao et al. [46], who utilize the Fisher Information Matrix (FIM) as a metric tensor to characterize the robustness of deep learning models. Building on this idea, we define a novel FIM-variant matrix, $\mathbf{H}\in\mathbb{R}^{n\times n}$, to evaluate the vulnerability of $\mathrm{LM}$ to perturbations in its feature space, with $\nabla_{x}\mathcal{J}(x,\hat{x})$ representing partial derivative of $\mathcal{J}(x,\hat{x})$ with respect to $x$:
$\mathbf{H(x)}=\nabla_{x}\mathcal{J}(x,\hat{x})^{\top}\nabla_{x}\mathcal{J}(x,\hat{x}).$ (10)
@@
TABLE III: Experimental results for the TOFU dataset using LLaMA2-7B for Forget05 subset. PerMU outperforms the baselines with a relative improvement of up to 50.40% (78.67 $\rightarrow$39.02) in Forget Probability and 11.3% (56.27$\rightarrow$62.63) on Forget Truth Ratio. Meanwhile, PerMU maintains the highest model utility, effectively preserving retained knowledge.

### v3

@@ Abstract
In this paper, we investigate knowledge forgetting in large language models with a focus on its generalisation—ensuring that models forget not only specific training samples but also related implicit knowledge. To this end, we begin by identifying a broader unlearning scope that includes both target data and logically associated samples, including rephrased, subject-replaced, relation-reversed, and one-hop reasoned data. We then conduct a rigorous evaluation of 15 state-of-the-art methods across three datasets, revealing that unlearned models still recall paraphrased answers and retain target facts in their intermediate layers. This motivates us to take a preliminary step toward more generalised implicit knowledge forgetting by proposing PerMU—a novel probability perturbation-based unlearning paradigm. PerMU simulates adversarial unlearning samples to eliminate fact-related tokens from the logit distribution, collectively reducing the probabilities of all answer-associated tokens. Experiments are conducted on a diverse range of datasets, including TOFU, Harry Potter, ZsRE, WMDP, and MUSE, using models ranging from 1.3B to 13B in scale. The results demonstrate that PerMU delivers up to a 50.40% improvement in unlearning vanilla target data while maintaining a 40.73% boost in forgetting implicit knowledge. Our code can be found in https://github.com/MaybeLizzy/PERMU.
@@ 1 Introduction
To further substantiate this observation, our first contribution is a comprehensive evaluation of the unlearning capability of existing methods in forgetting neighboring implicit knowledge. The evaluation covers three data domains: two widely-used machine unlearning datasets, TOFU [18] and Harry Potter [19, 20], as well as a popular model editing dataset, ZsRE [21]. We conduct experiments across 15 existing methods on two language models of different scales, Phi-1.3B and LLaMA2-7B. Our empirical analyses reveal the following unique findings that have been overlooked by existing research:
@@
  • The 1st cause: Unlearned models tend to remember target facts in their middle layers during inference;
  • The 2nd cause: Unlearned models are still capable of recalling paraphrased answers during inference.
@@
Motivated by these findings, we highlight a critical challenge in LLM unlearning: enhancing generalisation in forgetting implicit knowledge. However, achieving this goal is not without challenges. One vanilla approach is to identify and label all relevant latent knowledge (e.g., rephrased versions and paraphrased answers). Yet, this process is prohibitively labour-intensive and impractical, motivating the development of a novel solution.
@@
Building on the analysis using MSM, PerMU injects random noise into the embeddings of the most sensitive tokens, disrupting the model’s ability to recall factual information. As a result, the top-ranked tokens in the next-token probability distribution shift away from fact-related terms and toward tokens driven primarily by grammatical structure or contextual cues. Since the probability distribution reflects a language model’s internal knowledge [22], directly adjusting the logit distribution to suppress fact-related information offers an intuitive and effective unlearning strategy. PerMU implements this by subtracting the original distribution from the perturbed one, then fine-tuning the model to minimize the distance to this residual distribution—ensuring that correct fact-related tokens are assigned lower probabilities, thereby achieving unlearning. As such, PerMU simultaneously addresses both causes of the generalisation challenge as follows:
  • Solving the 1st cause: PerMU introduces random noise into the subject token embeddings at the first layer, effectively preventing the model from retrieving or generating factual information across subsequent middle layers;
  • Solving the 2nd cause: PerMU substantially reduces the probabilities of rich, highly-ranked answers and answer-related tokens in the original distribution, accomplished by the perturbed distribution subtraction.
In sum, our contributions are twofold: (1) a comprehensive evaluation of the generalization capability of existing unlearning methods in forgetting implicit knowledge; and (2) an unlearning method based on probability perturbation, PerMU, that effectively prevents models from recalling associated facts. We evaluate PerMU across five data domains, including the WMDP [23] and MUSE [24] datasets, using models of varying scales ($\sim13B$). Comparative experiments show that PerMU achieves up to a 50.40% improvement in unlearning and a 40.73% enhancement in generalisation, all while maintaining high model utility and superior generation quality.
@@ 2 Related Work
We provide a brief overview of existing machine unlearning methods for LLMs, categorised into training-free and training-based approaches [6], along with the evaluation methods.
@@
Unlearning Evaluation. Several studies have examined LLM unlearning from different perspectives [29, 30, 31, 32, 33]. Specifically, Patil et al. [25] investigate the effectiveness of typical model editing methods in removing information from model weights. Hong et al. [34] use vocabulary projections to analyze concept vectors through parametric knowledge traces. Yao et al. [35] evaluate seven different unlearning methods on longer-context tasks across three source domains. Meanwhile, Shi et al. [24] assess six desirable properties of unlearned models across eight unlearning methods. Jia et al. [20] focus on the influence of second-order optimization on unlearning. Li et al. [23] examine malicious use scenarios in biosecurity, cybersecurity, and chemical security. Qiu et al. [36] evaluate four distinct unlearning methods for removing highly interconnected data. Du et al. [37] investigate the risk of knowledge leakage after unlearning. The survey by Liu et al. [17] highlights often-overlooked aspects of existing LLM unlearning research and introduces the concept of unlearning scope. In contrast, we offer a more precise and formalized definition of unlearning scope, marking a fundamental distinction from prior work. While Patil et al. [31] explore extraction attacks to recover “deleted” information from intermediate hidden states, their focus lies primarily on model editing techniques such as ROME [38]. In comparison, we provide a comprehensive evaluation of the generalisation ability of existing machine unlearning methods across multiple dimensions. Regarding the generalisation evaluation, Dang et al. [32] focuses on improving generalised model utility, by introducing small Gaussian noise to retain data. Unlike prior works, we aim to enhance the generalisation in unlearning implicit knowledge associated with the target fact.
@@ 2.1 Problem Definition
[…]
3 Experimental Setup
In this section, we provide a detailed overview of the experimental setups for the unlearning generalization evaluation, covering the datasets, evaluation metrics, baselines, and implementation details.
@@ 3.1 Dataset
Our evaluation spans three distinct data domains, including two widely-used machine unlearning datasets, TOFU [18] and Harry Potter (HP)  [19], as well as a popular model editing dataset, ZsRE [21].
@@ 3.2 Evaluation Metrics
[…]
@@
[…]
@@
3.3 Fifteen Evaluated Algorithms
We evaluate 15 efficient unlearning methods, including Gradient Ascent (GA[10], Direct Preference Optimization (DPO[41], Negative Preference Optimization (NPO[11], Task Vectors (TV[13], Who’s Harry Potter (WHP[19], ULD [14], RMU [23], ECO [9] and ICL [26, 8].
@@ 3.3 Fifteen Evaluated Algorithms
  • Representation Misdirection for Unlearning (RMU) [23] fine-tunes the model by perturbing activations on hazardous data while preserving activations on benign data to mitigate malicious use. The full loss is defined as a weighted combination of the forget loss and the retain loss:
    $L=L_{\text{forget}}+\alpha\cdot L_{\text{retain}}.$ (8)
    We exclude RMU from overall evaluation because its objective—unlearning an entire distribution of hazardous knowledge given limited samples—differs fundamentally from our focus on unlearning privacy- and copyright-related knowledge, which assumes full access to the forget set. This makes direct comparison potentially unfair. A discussion of RMU on WMDP dataset is provided in Sect. 7.3.
  • Embedding Corrupted Prompts (ECO) [9] is a training-free unlearning approach that employs a scope classifier to identify prompts requiring unlearning and uses zeroth-order optimization to learn corruption parameters, which modify prompt embeddings at inference time—thus achieving unlearning without updating the original model weights. Since ECO relies on a trained scope classifier and only the classifier checkpoints for the TOFU dataset have been released, we use the released checkpoints directly for fair comparison, without further evaluation on HP and ZsRE.
    ECO has several variants, including ECO-Rand Noise (with perturbation strengths ranging from 5 to 4096), ECO-Zero-Out, and ECO-Sign-Flip. Although ECO-RN (strength = 4096) and ECO-Sign-Flip achieve strong forgetting effects, they incur substantial costs: the fluency metric for these variants is significantly lower. This degradation stems from excessive noise, which impairs the model’s language understanding and generation, often producing repetitive or low-quality text. To ensure comparable utility and generation quality with other methods, we report the performance of ECO-Zero-Out.
@@
We combine GA, DPO, and NPO with these two regularizations using a retain weight $RW$, denoted as “+GDR” or “+KLR”. The total unlearning loss is given by $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{unlearn}}(\theta)+RW\cdot\mathcal{L}_{\mathrm{retain}}(\theta)$. Following previous work [24, 18], we set $RW=1$. The combination of GDR and KLR results in a total of fifteen unlearning methods. Retain refers to the retain model, fine-tuned exclusively on the retain set without exposure to any forget data, and is considered an upper bound.
@@ 3.4 Implementation Details
The evaluation spans two models, Phi-1.3B and LLaMA2-7B. When tested on TOFU, we use the checkpoints of the pre-trained target model from the TOFU Leaderboard [^1] https://huggingface.co/spaces/locuslab/tofu_leaderboard. For the Harry Potter and ZsRE datasets, we first fine-tune the model on the respective dataset before applying unlearning. The fine-tuning settings are as follows: learning rate of 3e-5, 10 epochs, batch size of 8, with a gradient accumulation step of 4. For Task Vector and WHP, to obtain the reinforced model for unlearning, we fine-tune the target model for 10 epochs using the same learning rate and batch size. For ULD, we obtain the assistant model by fine-tuning the target model using the default settings provided by [14]. For the unlearning process, the unlearning batch size is set to 8, with a gradient accumulation step of 4. The process is conducted over 5 epochs, using a default learning rate of 2e-5. Since different learning rates can result in varying trade-offs between forgetting and retention, we slightly adjust the learning rate for each method to ensure comparable levels of model utility. To ensure fairness, all other unlearning hyper-parameters follow the default settings for each respective unlearning algorithm. All results are averaged over three runs. We use one A100 GPU with 80 GB of RAM. Note that during fine-tuning and unlearning on LLaMA2-7B, we update all 7B model parameters.
4 From Data to Insights: Analysis
In this section, we present a comprehensive analysis on implicit knowledge unlearning and observe that existing unlearning methods exhibit limited generalization ability. A detailed explanation is provided below.
@@ 5 Proposed Approach: PerMU
To achieve more generalized unlearning, simply reversing the gradient descent loss on forget samples is insufficient to fully erase implicit knowledge. Recent studies reveal that a language model’s internal knowledge is largely encoded in its probability distribution [22], which has inspired state-of-the-art approaches such as ULD [14]. These methods suppress fact-related information by modifying the logit distribution: an assistant model is fine-tuned to overfit ground truth, producing sharp logits that are then subtracted from the original logits to reduce the probability of correct responses. However, this fine-tuning process focuses narrowly on ground truth while overlooking paraphrased answers, leading to inference-time risks of generating semantically equivalent alternatives. In addition, these methods rely heavily on the quality of the assistant model’s training, which limits their scalability and practicality.
Motivated by these limitations, we propose PerMU, a more robust and efficient logit-adjustment method for achieving generalised unlearning. Unlike prior approaches that rely on assistant models, PerMU constructs adversarial unlearning samples to jointly suppress all answer-related tokens in the logit space. This is accomplished through computationally efficient perturbations that yield a modified distribution, which is then used for logit subtraction and matching. To determine which tokens should be perturbed, we introduce a theoretically grounded model-sensitivity metric, MSM, and present the complete unlearning pipeline in the following subsection.
@@ 5.1 Model Sensitivity Metric
[…]
@@
[…]